OpenStudy (anonymous):

How do you find vertical tangents and cusps? f(x)=(x+4)^-4/7 at c=4 f(x)=7(x-2)^3/7 at c=2

OpenStudy (anonymous):

Specifically for these problems (a) f(x)=(x+4)−4/7 at c=−4. (b) f(x)=7(x−2)3/7 at c=2. I have the answers, but I don't know how to approach the method.

zepdrix (zepdrix):

So like, you know how you set your first derivative equal to zero.... and solving for x gives you `horizontal asymptotes` usually, yes?

zepdrix (zepdrix):

Where your first derivative is undefined is usually where you're running into these vertical asymptotes or cusps.

zepdrix (zepdrix):

Do you understand how to find f'(x) for the first problem?

zepdrix (zepdrix):

This first question seems poorly written. The function doesn't even exist at x=-4, how could have a vertical tangent there? Hmm that's strange :(

OpenStudy (jhannybean):

If the limits approach two different values, it usually means there is a cusp. :P

zepdrix (zepdrix):

The second one makes sense though :)\[\Large\rm f(x)=7(x-2)^{3/7}\]Then,\[\Large\rm f'(x)=\frac{3}{(x-2)^{4/7}}\]Looking for tangents,\[\Large\rm 0=\frac{3}{(x-2)^{4/7}}\]We can see that the first derivative doesn't exist at (x-2)^{4/7}=0 (because we can't divide by 0 in the land of math). Solving for x gives us x=2.

zepdrix (zepdrix):

So what's your question exactly SizzlinBeef? 0_o Just the whole process confusing?